# CALM-PDE: Continuous and Adaptive Convolutions for Latent Space Modeling of Time-dependent PDEs

### v1

@@ Appendix I Additional Details on the Baselines
The Fourier Neural Operator is a neural operator based on Fast Fourier Transforms (FFTs). We use the implementation of FNO from PDEBench (Takamoto et al., 2022) and use the model in an autoregressive fashion as proposed by Li et al. (2020a) as “FNO with RNN structure”. However, we use a curriculum strategy that combines autoregressive training with teacher-forcing training by slowly increasing the rollout length of the autoregressive rollout (Li et al., 2023a) and by doing a teacher-forcing prediction with the remaining timesteps of the trajectory (Takamoto et al., 2023). The strategy improves the performance compared to full autoregressive training. We use the Geo-FNO (Li et al., 2023c), a variant of the FNO that supports irregularly sampled domains, for the irregularly sampled spatial domains experiments. The Geo-FNO also uses the curriculum training strategy for improved training.
@@
The errors of the autoregressive models (FNO, Geo-FNO, and PIT) can be reduced by using a curriculum strategy, that slowly increases the rollout length (Li et al., 2023a) and performs teacher-forcing training with the remaining timesteps (Takamoto et al., 2023), compared to fully autoregressive training. Table 4 shows that the curriculum strategy performs significantly better compared to fully autoregressive training. Thus, we train the autoregressive models with the curriculum strategy.
@@
We train the models with a relative L2 loss (Li et al., 2020a, 2023a, 2023b; Chen & Wu, 2024), which offers the advantage of treating channels with small magnitudes on par with those having large magnitudes, ensuring balanced weighting, while also providing feedback when the PDE solution decays (i.e., the magnitudes decay) over time. Let $\boldsymbol{Y}\in\mathbb{R}^{N_{t}\times N\times N_{c}}$ be the ground truth and $\boldsymbol{\hat{Y}}\in\mathbb{R}^{N_{t}\times N\times c}$ the model’s prediction. $N_{t}$ is the number of timesteps or trajectory length, $N$ the number of spatial points (e.g., $4096=64\cdot 64$ for a resolution of $64\times 64$ for 2D), and $N_{c}$ the number of channels. Then, the loss is defined as
@@
[…]
@@
Table 5 shows the hyperparameters for FNO and Geo-FNO used in the experiments. We adopt comparable hyperparameters to those in Li et al. (2020a) for the 1D Burgers’ equation and the 2D Navier-Stokes equations. For the 3D Navier-Stokes equations, we follow the hyperparameters outlined in Takamoto et al. (2022), with the exception of reducing the batch size from 5 to 2. For Geo-FNO, we utilize the hyperparameters proposed by Li et al. (2023c) which prove to be effective.
@@
Table 6 shows the hyperparameters for OFormer used in the experiments. We employ similar hyperparameters for the 1D Burgers’ equation and adopt the same parameters for the 2D Navier-Stokes equations as presented in Li et al. (2023a). For the 3D Navier-Stokes equations, we reduce the embedding dimensions to ensure the model fits on an A100 GPU. Increasing the model for 3D Navier-Stokes is infeasible due to excessive GPU memory usage and memory constraints in the experiments. For the 2D airfoil dataset, we utilize the hyperparameters specified in Li et al. (2023a) and adopt them for the 2D cylinder dataset.
@@
Encoder Embedding Dimension 96 96 64 128 128
Encoder Out Dimension 96 192 128 128 128
@@
Decoder Embedding Dimension 96 384 256 128 128
@@
Table 7 shows the hyperparameters for PIT used in the experiments. We adopt the hyperparameters introduced in Chen & Wu (2024) for the 1D Burgers’ equation, with the modification of increasing the batch size from 8 to 64. For the 2D Navier-Stokes equations, we retain the same hyperparameters as specified in Chen & Wu (2024). To train the 3D Navier-Stokes model on an A100 GPU, we reduce the latent mesh, quantiles, and the number of heads. Increasing the model for 3D Navier-Stokes is impractical due to GPU memory limitations encountered during the experiments. For the 2D airfoil and 2D cylinder datasets, we apply the hyperparameters used for the 2D Navier-Stokes equations which prove to be effective. The latent mesh for the 2D airfoil dataset is sampled from the training mesh (cf. mesh prior initialization of CALM-PDE) while the model employs a uniform latent mesh for the 2D cylinder dataset.
@@
Table 8 shows the hyperparameters for LNO used in the experiments. For the 1D Burgers’ equation, the listed hyperparameters prove to be effective. Compared to the hyperparameters introduced in Wang & Wang (2024) for the 2D Navier-Stokes equations, we decrease the embedding dimension, number of modes, and number of projector layers to better match the other models in terms of the number of parameters and required compute. For 3D Navier-Stokes, the embedding dimension is reduced further to train the model on an A100 GPU. Increasing the model is infeasible due to memory constraints in the experiments. We adopt the hyperparameters from 2D Navier-Stokes for the 2D airfoil dataset which prove to be effective. For the 2D cylinder dataset, we use similar hyperparameters as for the 2D cylinder dataset but reduce the embedding dimension from 192 to 176.
@@
Table 9 shows the hyperparameters for CALM-PDE used in the experiments. The notation [16, 32, 64] for the channels means that the first layer has 16 channels, the 2nd 32 channels, and the 3rd 64 channels. Similarly, for different hyperparameters in the table.
@@
Appendix K Model Analysis
@@
We visualize the learned positions of the query points of CALM-PDE. We conduct the experiment on models trained on the 2D Navier-Stokes equation, 2D Euler equation airfoil, and 2D Navier-Stokes cylinder datasets. Figure 7 shows the latent positions for a regular mesh where CALM-PDE samples more regularly. Figure 8 shows the input and output positions as well as learned query positions for the airfoil dataset. The results show that CALM-PDE samples more densely in important regions such as the boundary of the airfoil (see query points of decoder layer 2). A similar, but less emphasized effect can be observed for the encoder layer 1. Encoder layer 2 and decoder layer 1 do not show such an effect. Figure 9 shows the positions for the cylinder dataset. Similar to the airfoil dataset, CALM-PDE samples more densely in the regions where the cylinders are located. This effect can be observed in encoder layer 1 and decoder layer 2. Thus, the learned positions correspond to intuitively important regions and learnable query points help the model to have a higher information density in these important regions.
@@
Next, we investigate the information content of the query points. We are mainly interested in whether query points represent local or global information. Local information means that each query point represents the solution of the nearby surroundings while global information means that a single query point influences the global spatial domain. Due to the constrained receptive field, we assume that each token only represents local information. We add Gaussian noise to a token and investigate how the solution is affected to validate the hypothesis. Figure 10 shows that only the surrounding of the token is affected which matches our hypothesis that each token represents local information.
@@
Additionally, we visualize the learned kernel weights. Figure 11 shows the absolute values and Figure 12 the weights with the sign for the first encoder layer of CALM-PDE trained on the 2D Navier-Stokes dataset. The filters have similarities to the filters learned by a discrete CNN and some filters are similar to an edge detection filter (e.g., filter 4 in Figure 12). Figure 13 and Figure 14 show the learned filters for the second encoder layer of CALM-PDE. We only visualize the 64 filters for the first input channel. The filters are less regular and have similarities to attention weights in Transformers. It is important to note that filters complement each other by having one filter that focuses on a subset of tokens and another filter that focuses on the remaining subset of tokens (e.g., filters 31 and 38 in Figure 13).
@@
Appendix L Ablation Study
We conduct an ablation study to investigate the effect of (i) learnable query positions, (ii) kernel modulation, and (iii) distance weighting. Furthermore, we replace the proposed kernel function with a function that outputs weights only based on the Euclidean distance. We test the modified models on the 2D Euler equation airfoil dataset and the 2D Navier-Stokes dataset.
L.1 Learnable Query Positions
In the first experiment, we investigate the effect of learnable query positions. We conduct the experiment on the 2D Euler equation airfoil dataset with an irregularly sampled spatial domain. We examine three different configurations of CALM-PDE with (a) learnable query points and mesh prior initialization enabled (i.e., we initially sample the query points from the mesh and allow the model to change their positions), (b) no learnable query points but with mesh prior (i.e., we sample from the mesh to obtain the query points and their positions are fix), and (c) no learnable query points and no mesh prior (i.e., the query points are randomly sampled from the entire spatial domain and fixed). Table 11 shows that the model with learnable query points and mesh prior achieves the lowest errors. Soley sampling from the mesh to obtain the query points (mesh prior and no learnable query points) yields a significantly higher error. Using random query points with fixed positions results in the highest errors which is intuitive since the model does not necessarily have query points at important regions and cannot move query points to these regions. Thus, it is forced to learn the underlying mesh only with the kernel function.
@@
Model Configuration Relative L2 Error ($\downarrow$)
Learnable Query Points Mesh Prior 2D Euler Airfoil
@@
L.2 Kernel Modulation
In this experiment, we disable the query modulation. We conduct the experiment on the 2D Navier-Stokes dataset with a regularly sampled spatial domain and the 2D Euler airfoil dataset with an irregularly sampled spatial domain. We compare the CALM-PDE model with (a) enabled kernel modulation against a model (b) without kernel modulation. Table 12 shows the relative L2 errors for both model configurations. The proposed kernel modulation, which allows each query point to have a different kernel function, results in a lower error than the models without the kernel modulation.
@@
Kernel Modulation 2D Navier-Stokes $\nu=1e^{-4}$ 2D Euler Airfoil
@@
where $\mathtt{MLP}$ denotes the modulated two-layer multi-layer perceptron and $\mathtt{RFF}$ random Fourier features. Table 13 shows that the models without distance weighting and without a normalization term diverge during training. Including a normalization term of $1/M$ stabilizes the training but still yields a higher error than the model with distance weighting. Thus, distance weighting, which emphasizes input points closer to the query point and serves as a normalization, is crucial for reducing the error.
@@
Model Configuration Relative L2 Error ($\downarrow$)
Distance Weighting Normalization 2D Navier-Stokes $\nu=1e^{-4}$
@@
L.4 Distance-based Kernel Function
@@
where $T_{i,o}$ represents a learnable temperature parameter and $\mathtt{RF}(\alpha)$ the receptive field for the query point $\alpha$. We compare the CALM-PDE model with (a) the proposed CALM-PDE kernel function against a model with (b) the distance-based kernel function. Table 14 shows that the kernel function proposed by CALM-PDE achieves a lower error than a distance-based kernel function.
@@
Model Configuration Relative L2 Error ($\downarrow$)
Kernel Function 2D Navier-Stokes $\nu=1e^{-4}$
@@
Appendix M Additional Results
M.1 Quantitative Results
This section provides the full benchmark results with the standard deviations of multiple runs, which we omit for a more compact representation in the tables in the main paper. Table 15 shows the errors for the experiments with regular meshes and Table 16 for the irregular meshes.
@@
2D Euler Airfoil 2D Navier-Stokes Cylinder
@@
  • 1D Burgers’ Equation: Figure 15
  • 2D Navier-Stokes with Parameters $\nu=1e^{-4}$ and $\nu=1e^{-5}$: Figure 16
  • 2D Euler Equation with Airfoil Geometry: Figure 17, Figure 18, Figure 20, Figure 19
  • 2D Incompressible Navier-Stokes with Cylinder Geometries: Figure 21, Figure 22, Figure 23

### v2

@@ Appendix I Additional Details on the Baselines
The Fourier Neural Operator is a neural operator based on Fast Fourier Transforms (FFTs). We use the implementation of FNO from PDEBench (Takamoto et al., 2022) and use the model in an autoregressive fashion as proposed by Li et al. (2021b) as “FNO with RNN structure”. However, we use a curriculum strategy that combines autoregressive training with teacher-forcing training by slowly increasing the rollout length of the autoregressive rollout (Li et al., 2023a) and by doing a teacher-forcing prediction with the remaining timesteps of the trajectory (Takamoto et al., 2023). The strategy improves the performance compared to full autoregressive training. We use the Geo-FNO (Li et al., 2023c), a variant of the FNO that supports irregularly sampled domains, for the irregularly sampled spatial domains experiments. The Geo-FNO also uses the curriculum training strategy for improved training.
Factorized Fourier Neural Operator (F-FNO).
The Factorized Fourier Neural Operator (Tran et al., 2023) is an improved version of the Fourier Neural Operator that uses separable spectral convolution layers, improved residual connections, and pointwise MLPs. Separable spectral convolution layers factorize the Fourier transforms over the spatial dimension, which significantly decreases the number of model parameters. The improved residual connections, which are applied after the non-linearity, and the pointwise MLPs, akin to the pointwise MLPs in Transformers, further improve the performance. We use the original implementation of F-FNO and train the model in a similar fashion as FNO and Geo-FNO (i.e., autoregressive rollout with curriculum strategy).
@@
Attentive Reduced Order Model with Attention (AROMA).
AROMA (Serrano et al., 2024) is a reduced-order model that compresses the input solution into a fixed latent representation by applying cross-attention in a Perceiver-like fashion (Jaegle et al., 2021). A Transformer computes the latent representation for the next timestep with denoising diffusion, and a decoder maps the latent representation back to the physical space. Similar to LNO, AROMA directly supports irregularly sampled spatial domains. The authors use a two-stage training procedure that consists of an autoencoder training that trains the encoder-decoder model with a self-reconstruction loss and a dynamics training that trains the denoising diffusion model to predict the latent representation of the next timestep. We use their original implementation and keep the training strategy.
@@
The errors of the autoregressive models (FNO, Geo-FNO, and PIT) can be reduced by using a curriculum strategy, that slowly increases the rollout length (Li et al., 2023a) and performs teacher-forcing training with the remaining timesteps (Takamoto et al., 2023), compared to fully autoregressive training. Table˜4 shows that the curriculum strategy performs significantly better compared to fully autoregressive training. Thus, we train the autoregressive models with the curriculum strategy.
@@
We train the models with a relative L2 loss (Li et al., 2021b, 2023a, 2023b; Chen & Wu, 2024), which offers the advantage of treating channels with small magnitudes on par with those having large magnitudes, ensuring balanced weighting, while also providing feedback when the PDE solution decays (i.e., the magnitudes decay) over time. Let $\boldsymbol{Y}\in\mathbb{R}^{N_{t}\times N\times N_{c}}$ be the ground truth and $\boldsymbol{\hat{Y}}\in\mathbb{R}^{N_{t}\times N\times c}$ the model’s prediction. $N_{t}$ is the number of timesteps or trajectory length, $N$ the number of spatial points (e.g., $4096=64\cdot 64$ for a resolution of $64\times 64$ for 2D), and $N_{c}$ the number of channels. Then, the loss is defined as
@@
where $\left\lVert\cdot\right\rVert$ denotes the L2 norm.
@@
Table˜5 shows the hyperparameters for FNO and Geo-FNO used in the experiments. We adopt comparable hyperparameters to those in Li et al. (2021b) for the 1D Burgers’ equation and the 2D Navier-Stokes equations. For the 3D Navier-Stokes equations, we follow the hyperparameters outlined in Takamoto et al. (2022), with the exception of reducing the batch size from 5 to 2. For Geo-FNO, we utilize the hyperparameters proposed by Li et al. (2023c) which prove to be effective.
@@
Factorized Fourier Neural Operator (F-FNO).
Table˜6 summarizes the hyperparameters used for F-FNO in our experiments. For the 1D Burgers’ equation, we adopt hyperparameters similar to those in Li et al. (2021b), which prove to be effective. In the case of the 2D Navier-Stokes equations, we follow the configuration from Tran et al. (2023), but reduce the number of modes and the width to align with FNO, a necessary adjustment due to GPU memory limitations encountered during training. The 3D Navier-Stokes configuration also mirrors FNO’s hyperparameters. For the airfoil and cylinder datasets, we adopt the hyperparameters proposed by Li et al. (2023c).
Table 6: Hyperparameters for F-FNO used in the experiments.
Parameter PDE
1D Burgers’ 2D Navier-Stokes 3D Navier-Stokes 2D Airfoil 2D Cylinder
Width 64 24 20 24 24
Modes 16 14 12 14 14
Layers 4 4 4 6 6
Learning Rate 1e-3
Batch Size 32 64 2 8 8
Epochs 500
Parameters 665,281 151,361 131,913 625,270 625,229
@@
Table˜7 shows the hyperparameters for OFormer used in the experiments. We employ similar hyperparameters for the 1D Burgers’ equation and adopt the same parameters for the 2D Navier-Stokes equations as presented in Li et al. (2023a). For the 3D Navier-Stokes equations, we reduce the embedding dimensions to ensure the model fits on an A100 GPU. Increasing the model for 3D Navier-Stokes is infeasible due to excessive GPU memory usage and memory constraints in the experiments. For the 2D airfoil dataset, we utilize the hyperparameters specified in Li et al. (2023a) and adopt them for the 2D cylinder dataset.
@@
Encoder
Embedding Dimension
96 96 64 128 128
Encoder
Out Dimension
96 192 128 128 128
@@
Decoder
Embedding Dimension
96 384 256 128 128
@@
Table˜8 shows the hyperparameters for PIT used in the experiments. We adopt the hyperparameters introduced in Chen & Wu (2024) for the 1D Burgers’ equation, with the modification of increasing the batch size from 8 to 64. For the 2D Navier-Stokes equations, we retain the same hyperparameters as specified in Chen & Wu (2024). To train the 3D Navier-Stokes model on an A100 GPU, we reduce the latent mesh, quantiles, and the number of heads. Increasing the model for 3D Navier-Stokes is impractical due to GPU memory limitations encountered during the experiments. For the 2D airfoil and 2D cylinder datasets, we apply the hyperparameters used for the 2D Navier-Stokes equations which prove to be effective. The latent mesh for the 2D airfoil dataset is sampled from the training mesh (cf. mesh prior initialization of CALM-PDE) while the model employs a uniform latent mesh for the 2D cylinder dataset.
@@
Table˜9 shows the hyperparameters for LNO used in the experiments. For the 1D Burgers’ equation, the listed hyperparameters prove to be effective. Compared to the hyperparameters introduced in Wang & Wang (2024) for the 2D Navier-Stokes equations, we decrease the embedding dimension, number of modes, and number of projector layers to better match the other models in terms of the number of parameters and required compute. For 3D Navier-Stokes, the embedding dimension is reduced further to train the model on an A100 GPU. Increasing the model is infeasible due to memory constraints in the experiments. We adopt the hyperparameters from 2D Navier-Stokes for the 2D airfoil dataset which prove to be effective. For the 2D cylinder dataset, we use similar hyperparameters as for the 2D cylinder dataset but reduce the embedding dimension from 192 to 176.
@@
AROMA.
Table˜10 summarizes the hyperparameters used for AROMA in our experiments. We adopt the hyperparameters reported in Serrano et al. (2024), but reduce the number of epochs to finish each training within 24 hours.
Table 10: Hyperparameters for AROMA used in the experiments.
Parameter PDE
1D Burgers’ 2D Navier-Stokes 3D Navier-Stokes 2D Airfoil 2D Cylinder
Encoder-Decoder
Hidden Dimension 128
Number of Latents 32 256 256 64 64
Latent Dimension 8 16 16 16 16
Depth 3
Encode Geometry
Learning Rate 1e-3
Batch Size 128 10 4 32 64
Epochs 1500 1000 500 5000 10000
Dynamics Model
Hidden Dimension 128
Depth 4
Minimum Noise 1e-2 1e-3 1e-3 1e-6 1e-3
Denoising Steps 3
Learning Rate 1e-3
Batch Size 512 64 4 32 128
Epochs 1000
Parameters 665,281 1,845,505 1,852,773 1,950,148 1,949,891
@@
Table˜11 shows the hyperparameters for CALM-PDE used in the experiments. The notation [16, 32, 64] for the channels means that the first layer has 16 channels, the 2nd 32 channels, and the 3rd 64 channels. Similarly, for different hyperparameters in the table. The dimensionality of the latent space is defined by the number of latent variables or query points and their channel dimension. We determine the dimensionality of the latent space based on the principle that the number of query points, receptive field, and channel dimension must cover the entire spatial domain. Similar to discrete CNNs, a small receptive field with a large stride (e.g., only a few query points) misses information, while a large receptive field with a small stride could have an averaging effect. With this principle, we obtain the number of query points (e.g., 1024, 256, 16 for the encoder layers), which we use for all 2D experiments. Thus, a fixed number of query points works across different problems.
@@
Appendix K Hyperparameter Study
This section provides additional information about the hyperparameters of CALM-PDE and their robustness. In particular, we show the similarities of the hyperparameters of CALM-PDE and discrete CNNs, and demonstrate the robustness of the softmax temperature.
Comparison to discrete CNNs.
CALM-PDE introduces the softmax temperature, receptive field radius (percentile), channel dimension, and number of query points as hyperparameters. The receptive field radius (percentiles) is equivalent to the kernel size in discrete CNNs, and the number of query points to downsampling (e.g., stride and pooling) in CNNs. The channel dimension is equivalent to the channel dimension in CNNs. Alternatively, the number of query points can be compared to the number of latent query tokens used in other models, such as AROMA (Serrano et al., 2024) or LNO (Wang & Wang, 2024). Consequently, CALM-PDE maintains a conventional hyperparameter footprint without introducing excessive complexity.
Hyperparameter Selection and Robustness.
Selecting hyperparameters for CALM-PDE does not require an extensive hyperparameter search. As shown in Table˜11, we use the same temperatures, receptive field sizes (percentiles), and number of query points for all 2D experiments, which demonstrates that the hyperparameters are robust and work across different problems.
Hyperparameter Study of Softmax Temperature.
The softmax temperature is a new hyperparameter that controls the distance weighting within the receptive field. We also use the same temperatures across all experiments, demonstrating its robustness. Additionally, we perform a hyperparameter study by sweeping the temperature from a high value in model (a) to a low value in model (f). As shown in Table˜12, temperatures in the range of $T\in[0.1,5.0]$ result in only minor changes in test error, confirming the robustness of the model to this parameter. All experiments are conducted on the 2D Navier-Stokes dataset with $\nu=1e^{-4}$.
Table 12: Relative L2 errors of CALM-PDE models with different temperatures trained and tested on the 2D Navier-Stokes dataset.
Model Configuration
Relative L2 Error ($\downarrow$)
Temperature T in Encoder and Decoder
2D Navier-Stokes
$\nu=1e^{-4}$
(a) 10 0.0398
(b) 5 0.0336
(c) 2 0.0316
(d) 1 0.0304
(e) 0.1 0.0341
(f) 0.01 0.1640
Appendix L Model Analysis
@@
We visualize the learned positions of the query points of CALM-PDE. We conduct the experiment on models trained on the 2D Navier-Stokes equation, 2D Euler equation airfoil, and 2D Navier-Stokes cylinder datasets. Figure˜7 shows the latent positions for a regular mesh where CALM-PDE samples more regularly. Figure˜8 shows the input and output positions as well as learned query positions for the airfoil dataset. The results show that CALM-PDE samples more densely in important regions such as the boundary of the airfoil (see query points of decoder layer 2). A similar, but less emphasized effect can be observed for the encoder layer 1. Encoder layer 2 and decoder layer 1 do not show such an effect. Figure˜9 shows the positions for the cylinder dataset. Similar to the airfoil dataset, CALM-PDE samples more densely in the regions where the cylinders are located. This effect can be observed in encoder layer 1 and decoder layer 2. Thus, the learned positions correspond to intuitively important regions and learnable query points help the model to have a higher information density in these important regions.
@@
Next, we investigate the information content of the query points. We are mainly interested in whether query points represent local or global information. Local information means that each query point represents the solution of the nearby surroundings while global information means that a single query point influences the global spatial domain. Due to the constrained receptive field, we assume that each token only represents local information. We add Gaussian noise to a token and investigate how the solution is affected to validate the hypothesis. Figure˜10 shows that only the surrounding of the token is affected which matches our hypothesis that each token represents local information.
@@
Additionally, we visualize the learned kernel weights. Figure˜11 shows the absolute values and Figure˜12 the weights with the sign for the first encoder layer of CALM-PDE trained on the 2D Navier-Stokes dataset. The filters have similarities to the filters learned by a discrete CNN and some filters are similar to an edge detection filter (e.g., filter 4 in Figure˜12). Figure˜13 and Figure˜14 show the learned filters for the second encoder layer of CALM-PDE. We only visualize the 64 filters for the first input channel. The filters are less regular and have similarities to attention weights in Transformers. It is important to note that filters complement each other by having one filter that focuses on a subset of tokens and another filter that focuses on the remaining subset of tokens (e.g., filters 31 and 38 in Figure˜13).
@@
Appendix M Ablation Study
We conduct an ablation study to investigate the effect of (i) learnable query positions, (ii) kernel modulation, and (iii) distance weighting. Furthermore, we replace the proposed kernel function with a function that outputs weights only based on the Euclidean distance. We also investigate the influence of different latent dimensions on the model’s performance. We test the modified models on the 2D Euler equation airfoil dataset and the 2D Navier-Stokes dataset.
M.1 Learnable Query Positions
In the first experiment, we investigate the effect of learnable query positions. We conduct the experiment on the 2D Euler equation airfoil dataset with an irregularly sampled spatial domain. We examine three different configurations of CALM-PDE with (a) learnable query points and mesh prior initialization enabled (i.e., we initially sample the query points from the mesh and allow the model to change their positions), (b) no learnable query points but with mesh prior (i.e., we sample from the mesh to obtain the query points and their positions are fix), and (c) no learnable query points and no mesh prior (i.e., the query points are randomly sampled from the entire spatial domain and fixed). Table˜14 shows that the model with learnable query points and mesh prior achieves the lowest errors. Soley sampling from the mesh to obtain the query points (mesh prior and no learnable query points) yields a significantly higher error. Using random query points with fixed positions results in the highest errors which is intuitive since the model does not necessarily have query points at important regions and cannot move query points to these regions. Thus, it is forced to learn the underlying mesh only with the kernel function.
@@
Model Configuration
Relative L2 Error ($\downarrow$)
Learnable Query Points Mesh Prior
2D Euler Airfoil
@@
M.2 Kernel Modulation
In this experiment, we disable the query modulation. We conduct the experiment on the 2D Navier-Stokes dataset with a regularly sampled spatial domain and the 2D Euler airfoil dataset with an irregularly sampled spatial domain. We compare the CALM-PDE model with (a) enabled kernel modulation against a model (b) without kernel modulation. Table˜15 shows the relative L2 errors for both model configurations. The proposed kernel modulation, which allows each query point to have a different kernel function, results in a lower error than the models without the kernel modulation.
@@
Kernel Modulation
2D Navier-Stokes
$\nu=1e^{-4}$
2D Euler Airfoil
@@
where $\mathtt{MLP}$ denotes the modulated two-layer multi-layer perceptron and $\mathtt{RFF}$ random Fourier features. Table˜16 shows that the models without distance weighting and without a normalization term diverge during training. Including a normalization term of $1/M$ stabilizes the training but still yields a higher error than the model with distance weighting. Thus, distance weighting, which emphasizes input points closer to the query point and serves as a normalization, is crucial for reducing the error.
@@
Model Configuration
Relative L2 Error ($\downarrow$)
Distance Weighting Normalization
2D Navier-Stokes
$\nu=1e^{-4}$
@@
M.4 Distance-based Kernel Function
@@
where $T_{i,o}$ represents a learnable temperature parameter and $\mathtt{RF}(\alpha)$ the receptive field for the query point $\alpha$. We compare the CALM-PDE model with (a) the proposed CALM-PDE kernel function against a model with (b) the distance-based kernel function. Table˜17 shows that the kernel function proposed by CALM-PDE achieves a lower error than a distance-based kernel function.
@@
Model Configuration
Relative L2 Error ($\downarrow$)
Kernel Function
2D Navier-Stokes
$\nu=1e^{-4}$
@@
M.5 Latent Dimension
We investigate the effect of different latent dimensions by increasing the number of query points. Starting from model (a), which uses a small number of query points, we incrementally scale up to model (e), characterized by a large number of query points. As shown in Table˜18, increasing the latent dimension initially leads to a reduction in test error. However, beyond a certain threshold, further increases of the latent dimension result in an increasing test error, suggesting overfitting or diminishing returns. The table also reports training time, which consistently increases with the number of query points, reflecting the computational cost of a higher latent dimension. The notation [256, 64, 4] denotes 256 query points for the first layer, 64 query points in the 2nd layer, and 4 for the last layer.
Table 18: Relative L2 errors and training times of CALM-PDE models with different numbers of query points trained and tested on the 2D Navier-Stokes dataset.
Model Configuration
Relative L2 Error ($\downarrow$)
Training time ($\downarrow$)
Number of Query Points
2D Navier-Stokes
$\nu=1e^{-4}$
Encoder Decoder
(a) [256, 64, 4] [64, 256, -] 0.0533 4:18h
(b) [512, 128, 8] [128, 512, -] 0.0425 4:40h
(c) [1024, 256, 16] [256, 1024, -] 0.0304 5:25h
(d) [2048, 512, 32] [512, 2048, -] 0.0251 6:21h
(e) [4096, 1024, 64] [1024, 4096, -] 0.0267 8:19h
Appendix N Additional Results
N.1 Quantitative Results
This section provides the full benchmark results with the standard deviations of multiple runs, which we omit for a more compact representation in the tables in the main paper. Table˜19 shows the errors for the experiments with regular meshes and Table˜20 for the irregular meshes.
@@
1D Burgers’
$\nu=1e^{-3}$
2D Navier-Stokes
$\nu=1e^{-4}$
2D Navier-Stokes
$\nu=1e^{-5}$
3D Navier-Stokes
$\eta=\zeta=1e^{-8}$
@@
2D Euler Airfoil
2D Navier-Stokes Cylinder